OpenStudy (anonymous):

solve for g. 6g-23=12-g

OpenStudy (anonymous):

g=5

OpenStudy (anonymous):

first of all, group like terms... Send all the g terms to one side of the equal sign and send the constants to the other side of the equal sign..

OpenStudy (anonymous):

Thanks i got 5 to i just second guess my self

OpenStudy (anonymous):

yes...the answer is 5.

OpenStudy (unklerhaukus):

add g to both sides, and add 23 to both sides, then divide by seven

OpenStudy (anonymous):

solve for d. d2^ -11=5 can you help me on this one to?

OpenStudy (unklerhaukus):

is that \[d^2 -11=5\]

OpenStudy (anonymous):

yes.

OpenStudy (unklerhaukus):

add eleven to both sides, then take the square root of both sides, dont forget that \[\sqrt{x^2}=\pm x\]

OpenStudy (anonymous):

would it be 4

OpenStudy (anonymous):

idk the second number sorry im not good at this

OpenStudy (unklerhaukus):

well yes 4 is one solution , there is one more solution though

OpenStudy (anonymous):

will the second one be -4?

OpenStudy (unklerhaukus):

right! because \[(-4)^2−11=5\]\[16-11=5\]

OpenStudy (anonymous):

a. 4, –4 b.5, –5 c.5, 6 d.6, –6 these are the answer choices

OpenStudy (unklerhaukus):

well that good, because the two solutions you found form a pair of solution in one of those options

OpenStudy (anonymous):

Hmm so the answer would be A right ?

OpenStudy (unklerhaukus):

\[\checkmark\]